# Machine Learning Unveils the power law of Finite-Volume Energy Shifts

### v1

@@ Abstract
Finite-volume extrapolation is an important step for extracting physical observables from lattice calculations. However, it is a significant challenge for the system with long-range interactions. We employ symbolic regression to regress finite-volume extrapolation formula for both short-range and long-range interactions. The regressed formula still holds the exponential form with a factor $L^{n}$ in front of it. The power decreases with the decreasing range of the force. When the range of the force becomes sufficiently small, the power converges to $-1$, recovering the short-range formula as expected. Our work represents a significant advancement in leveraging machine learning to probe uncharted territories within particle physics.
@@
Introduction:   Lattice Quantum Chromodynamics, which discretizes space into four dimensional hypercubic lattice and perform the calculation in a given finite volume, is a foundational computational framework for studying strong interaction. It predicts not only the spectrum of conventional hadrons but also reveals the existence of exotic hadronic states. Especially, it provides precise calculations for near-threshold states recently, taking the double charm tetraquark as an example [1, 2, 3]. To extract physical observables from lattice calculations, finite-volume extrapolation is essential. However, the presence of long-range interactions presents significant challenges, as they invalidate key assumptions underlying standard approaches, most notably the Lüscher formula [4]. When light particles mediate forces between scattering states, conventional methods may fail, leading to uncontrolled systematic uncertainties. Addressing these issues is crucial for reliable lattice analyses, particularly in systems where long-range effects play a dominant role, such as one-pion exchange in nucleon-nucleon scattering and the effects of light-meson exchange in the $D^{*}D$ system [5]. Recent theoretical studies [6, 7, 8, 9] try to solve this problem by either modifying Lüscher formula numerically, but without explicit formula as simple as that for the short-range interaction [10], or performing quantization conditions using plane wave basis [11]. Moreover, understanding finite-volume effects in two-body systems with long-range interactions is a necessary step toward addressing similar challenges in three-body systems.
@@
Fortunately, deriving fundamental mathematical expressions directly from empirical data [12] is an advantage of symbolic regression [13], which has exhibited its power in physics and scientific research [14]. Unlike traditional regression, it autonomously uncovers interpretable laws by exploring vast expression spaces. This has enabled rediscovery of known physical laws and formulation of novel equations for complex systems [15] even if their underlying principles were elusive. Recent advances in symbolic regression enhance its effectiveness by incorporating physical constraints like dimensional consistency, improving search efficiency and model plausibility. For example, the AI Feynman algorithm, derived 100 equations from the Feynman Lectures on Physics [16]. Its ability to generate human-readable models accelerates discovery across various fields [17, 18]. For instance in high-energy physics, it has revealed fundamental distributions, such as the Tsallis distribution in hadron transverse momentum analysis [19]. Even, it has provided analytical expressions for key low-energy observables, including the Higgs mass, muon anomalous magnetic moment, and dark matter relic density [20].
@@
As the result, we employ symbolic regression to probe the finite-volume extrapolation formula for long-range interactions. Before symbolic regression, we need to prepare samples for training symbolic regression. For these samples, we borrow the Hadron Lattice Effective Field Theory (HLEFT) [21], which is analogous to the Nuclear Lattice Effective Field Theory (NLEFT), a powerful computational framework for ab initio nuclear structure calculations [22]. The approach implements a systematic discretization of nucleon-nucleon interactions on a three-dimensional cubic lattice with periodic boundary conditions. These quantum many-body systems are subsequently solved using advanced numerical techniques including the Lanczos diagonalization method and auxiliary field Monte Carlo simulations. NLEFT has demonstrated remarkable success across diverse domains of nuclear physics, particularly in characterizing nuclear ground states [23, 24, 25, 26, 27, 28, 29] and excited states [30, 31, 32, 33], elucidating nuclear intrinsic density distributions and clustering phenomena [34, 35, 36, 37, 38], analyzing nucleus-nucleus scattering processes [39, 40], and investigating both zero- and finite-temperature nuclear matter properties [41, 42, 43, 44]. The framework has also been effectively extended to study hypernuclear systems [45, 46, 47, 21]. However, the presence of finite-volume effects remains a significant challenge in extending NLEFT applications to critical nuclear systems such as the neutron-rich ${}^{6}\mathrm{He}$ isotope, where the delicate balance of nuclear binding requires exceptional precision. This is another motivation of this work. According to the statement above, our workflow is presented in Fig. 1.
Figure 1: Workflow of using SR, LEFT and formula.
@@
Samples generation:   The HLEFT is used to generate the model-independent samples of two identical boson systems. These samples are fed to Symbolic Regression (SR), which is based on genetic algorithms (GA) [48, 49]. GA allows computer to use fewer prior assumptions and more flexible search space to find final expression. Similar to the evolutionary processes in nature, GA employs mechanisms such as mutations and crossovers to generate new solutions. In this work, the famous PySR, which is a multi-population evolutionary with multiple evolutions performed asynchronously. It uses a classic evolutionary algorithm for every cycle and tournament selection for individual selection. The details of the setup in PySR can be found in Ref. [50].
@@
The challenge for finite volume extrapolation is for long-range interactions [6, 11, 7, 8, 9], especially when the range of force is comparable with the box size. Luckily, the extrapolation formula for the short-range potential can be obtained from Lüscher’s formula for $\kappa L\gg 1$ [10], with $\kappa$ the binding momentum and $L$ the box size. Firstly, we regress the extrapolation formula for the short-range potential as a benchmark for the applicability of PySR model. We work in non-relativistic framework and consider two identical spinless particles with the same mass $m=1969~{}\mathrm{MeV}$.   [^1] The polarization of non-zero spin particles would modify the finite-volume extrapolation formula, which will not be discussed in this manuscript. The short-range potential in coordinate space reads as
$V(\bm{r})=-C_{0}\delta^{3}(\bm{r}).$ (1)
Here, $\bm{r}$ is the relative distance between the two particles. $C_{0}$ is the strength of the potential for tuning the binding energy of the two-particle system. The minus sign means an attractive potential. From effective field theory (EFT) point of view, the short-range potential $V(\bm{r})$ needs a physical cutoff to isolate the long-range contribution from the short one. Usually, a single-particle regulator [51] $f(\bm{p}^{(\prime)}_{1},\bm{p}^{(\prime)}_{2})=\prod\limits_{i=1}^{2}g_{\Lambda}(\bm{p_{i}})g_{\Lambda}(\bm{p^{\prime}_{i}})$ is applied, where $g_{\Lambda}(\bm{p})=$exp$(-\bm{p}^{6}/2\Lambda^{6})$ is a soft cutoff function with $\bm{p_{i}}$ and $\bm{p^{\prime}_{i}}$ the momenta of the incoming and outgoing of the $i$th particle [21]. Thus, the Hamiltonian can be written as:
$H=\sum_{i=1}^{2}\frac{\bm{p_{i}}^{2}}{2m_{i}}+f(\bm{p}^{(\prime)}_{1},\bm{p}^{(\prime)}_{2})V(\bm{q}).$ (2)
@@
Here, $m_{i}$ is the mass of the $i$th particle. $\bm{q}=\bm{p}-\bm{p}^{\prime}$ is transferred momentum of the two particles and $\bm{p}$($\bm{p^{\prime}}$) represents relative incoming (outgoing) momentum between $\bm{p}_{1}$ ($\bm{p}^{\prime}_{1}$) and $\bm{p}_{2}$ ($\bm{p}^{\prime}_{2}$). The corresponding potential in momentum space $V(\bm{q})$ = $-C_{0}$ can be obtained by a Fourier transform from the expression in coordinate space.
@@
$H_{L}$ and $E_{L}$ are the Hamiltonian and the binding energy in a cubic box $L^{3}$. The solution of the Schrödinger equation (Eq. 3) can be written explicitly in a single-particle basis $|\psi\rangle=|n_{1},\cdots,n_{N}\rangle$ [42], where $n_{i}$ is an integer triplet specifying the lattice coordinates. The matrix exact diagonalization scheme uses the implicitly restarted Lanczos method to find the eigenvalues and eigenvectors [52] by SciPy [53] in Python. In order to study the finite volume extrapolation formula for short-range potential, we perform simulations on box size $L^{3}=10^{3},11^{3},\cdots,30^{3}~{}\mathrm{fm}^{3}$ cubic lattice with two identical particles. The potential strength $C_{0}$ is set to 1.5, 2.0, 2.5 and 3.0 $\mathrm{MeV}^{-2}$ to generate the samples. With this setup, we can extract the energy $E_{L}$ in finite volume by solving Eq. (3) for various cases. The details of the samples can be found in Ref. [50].
The Yukawa potential [6]
$V(\bm{r})=-C_{01}\delta^{3}(\bm{r})-C_{02}\frac{e^{-\mu r}}{r}.$ (4)
with $1/\mu$ reflecting the range of the force is used to describe the long-range potential. Here, $C_{01}$ and $C_{02}$ are the strengths of the short-range potential ($\delta$-potential) and the long-range potential (Yukawa potential). One notice that the short-range behavior of Yukawa potential is divergent and needs regularization, which is the reason why the first term exists. For simplicity, we set $C_{01}$ = $C_{02}$ and choose the values 0.03, 0.09, 0.12 and 0.21 MeV-2 to generate the samples. These values allow the binding energy of the two-body system with the long-range potential close to that with only the short-range potential. Only in this case, one can isolate the role of the long-range potential on the finite-volume extrapolation formula. For the short-range term, the single-particle regulator $f(\bm{p}^{(\prime)}_{1},\bm{p}^{(\prime)}_{2})$ is still applied. At the same time, for the long-range potential term, the regulator222 Here, $\Lambda$ is the cutoff to regularize the long-range potential, which is set equal to that for the short-range one above. $\hat{f}(\bm{q})$ = exp$(-(\bm{q}^{2}+\mu^{2})/\Lambda^{2})$ is applied [5]. The Hamilton $H$ in this case can be written as:
@@
Here, both the short-range potential $V_{S}(\bm{q})$ = $-C_{01}$ and the long-range potential $V_{L}(\bm{q})\sim\frac{1}{\bm{q}^{2}+\mu^{2}}$ can be transformed into their coordinate form by Fourier transforms. We perform simulations on box size $L^{3}=10^{3},11^{3},\cdots,30^{3}$ fm3 cubic lattice. The potential strength $C_{01}=C_{02}=0.03,~{}0.09,~{}0.12,~{}0.21~{}\mathrm{MeV}^{-2}$ to obtain binding energies similar to those for only the short-range potential. The details can be found in Ref. [50].
Symbolic regression:  We perform symbolic regression on the generated samples that characterize finite-volume effects. The PySR [54] model samples the space of analytic expressions defined by the sign of operators, input variables, and constant terms for minimization through genetic programming. The pool of operations are addition, subtraction, multiplication, division, exponential, logarithm and square, etc. The input variable is box size $L$. Unlike massive data used in deep learning, a few samples should be sufficient in PySR model. Model evolution progress employs several mechanisms, such as mutations and crossovers, to generate new expressions. We mutate over 50 iterations of 50 different population samples, with each population containing 35 individuals. There are two elements to measure the goodness of the output formula in PySR model, i.e. loss and score. The value of loss is used to measure how well the output formula describes the samples, which is defined as mean square error:
@@
Here, $E_{\mathrm{PySR}}(L)$ is the output formula of PySR model, $E_{L}({L})$ is the energy at box size $L^{3}$ calculated from LEFT. Score is used to estimate the form of the formula, which rewards minimal value of loss and penalizes the more complicated formula. It can be defined as:
@@
Here, $C$ is the complexity  [^3] The upper limit of the $C$ value can be set larger to allow higher order contribution as that in analytic analysis  [10]. (its values can be found in Ref. [50]), which is defined as the total numbers of operations, variables, and constants used in a formula. PySR model will consider both score, i.e. complexity and loss, to choose the best formula. The other parameters used in training process can be found in Ref. [50].
@@
The training process of SR is very similar to biological population inheritance. In initialization, PySR model randomly provides several kinds of evolutionary frameworks, which are different from each other and are used as a starting point for a given evolution. Mutations and crossovers are throughout each iteration. During each iteration, formulae will be produced and PySR model will choose a best formula into the next iteration according to the value of loss and score. The iterations will stop as long as the loss and complexity are up to the specified value set at the beginning. During this process, the PySR model will choose the best formula throughout iterations.
@@
Figure 2: The results of fitting the formula to samples. (a) is the case of $C_{0}=1.5$ MeV-2, (b) is the case of $C_{0}=2.0$ MeV-2, (c) is the case of $C_{0}=2.5$ MeV-2 and (d) is the case of $C_{0}$ = 3.0 MeV-2. Eqs. (A), (B) and (C) correspond to Eqs. (8A), (8B) and (8C), respectively. The box size $L$ ranges from 10 to 30 fm. The purple region and the yellow region indicate the fitted and predicted area by Eqs. (A), (B) and (C), respectively. Each graph has its own sub-graph to show more detailed of lines between fitted and predicted area.
@@
Results:  Based on the PySR model, we can perform a symbolic regression for the samples of the short-range attractive potential to find an expression that uniformly describes the finite-volume extrapolation. In the four cases of $C_{0}$ values, three formulae  [^4] For the case of $C_{0}=1.5~{}\mathrm{MeV}^{-2}$, PySR model gives the formula $E_{L}=C^{\prime}_{1}+C^{\prime}_{2}/L+C^{\prime}_{3}/L^{2}$. For the purpose of uniformity of the formula, when the ratio $C^{\prime}_{2}/C_{3}^{\prime}$ is small enough, the formula is approximated as $E_{L}=C_{1}+C_{2}e^{-C_{3}L}/L^{2}$. are accepted by PySR model, as follows,
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L^{2},$ (8A)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L,$ (8B)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}.$ (8C)
@@
In Fig. 2, we show the results of the three formulae fitting to the samples with the short-range potential, generated from LEFT. The purple area is the fitting region with the formulae derived from PySR model. The yellow area is the predicted one. For the shallow bound state, i.e. Fig. 2 (a), Eq. (8A) is the best choice. However, the volume size is not big enough to see the convergent behavior, making the regressed formula not universal. When binding energy increases to a few MeV, i.e. Fig. 2 (b) and (c), Eq. (8B) is better than Eq. (8A). When binding energy further increases to more than $10~{}\mathrm{MeV}$, i.e. Fig. 2 (d), all the three formulae are good choices because their loss values are much smaller than the other subfigures of Fig. 2. As the result, we conclude that Eq. (8B) is the best regressed uniform formula to describe the samples with a few MeV binding energy.
@@
The current well-established theoretical works [4, 55, 56, 57] are based on the Lüscher formula [10] for the short-range potential with $\kappa L\gg 1$ and $\kappa$ is the binding momentum of the two-body system. The energy shift of a shallow two-body bound state of identical bosons in finite volume is
@@
with $C^{\prime}$ is the free parameter. This formula is exactly the form of Eq. (8B) regressed by PySR. In addition, the fitted values of the parameter $C_{3}$ in Eq. (8B) for the last three cases of Fig. 2 equal to the binding momentum $\kappa$ numerically. This also demonstrates that we have successfully regressed the finite-volume extrapolation formula for short-range potential.
Figure 3: The results of Eq. (10A) (green short dashed curves), Eq. (10B) (red dashed curves) and Eq. (10C) (orange long dashed curves) fitting to the samples of the long-range potentials with $C_{01}=C_{02}=0.03~{}\mathrm{MeV}^{-2}$ (a), $C_{01}=C_{02}=0.09~{}\mathrm{MeV}^{-2}$ (b), $C_{01}=C_{02}=0.12~{}\mathrm{MeV}^{-2}$ (c), and $C_{01}=C_{02}=0.21~{}\mathrm{MeV}^{-2}$ (d). The box size $L$ ranges from $10\sim 30~{}\mathrm{fm}$. The purple region and the yellow region indicate the fitted and predicted area. Each graph has its own sub-graph to show more details around the boundary between fitted and predicted area. The black points are the samples generated from LEFT.
For the long-range potential, the strengths of the attractive potentials are set to obtain similar binding energies for the short-range potential. The range of the force parameter $\mu$ is set as $\mu=20~{}\mathrm{MeV}$ for these four cases. The other values of this parameter and their effect on the finite-volume extrapolation will be discussed afterwards. The following three formulae  [^5] For the case of $C_{01}=C_{02}=$ 0.03 MeV-2, PySR model gets the formula $E_{L}=C^{\prime}_{1}+C^{\prime}_{2}/L+C^{\prime}_{3}L$. For the purpose of a uniform form of the formula, when the ratio $C^{\prime}_{3}/C^{\prime}_{2}$ is small enough, the formula is approximated as $E_{L}=C_{1}+C_{2}e^{-C_{3}L}/L$.
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L,$ (10A)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L},$ (10B)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}L.$ (10C)
are accepted by PySR model. Fig. 3 shows the fitting results of the above three formulae to the energy with the long-range interaction. For the case with binding energy smaller than $1~{}\mathrm{MeV}$, i.e. Fig. 3(a), Eq. (10A) is preferred. However the binding energy is too small to make the energy convergent in box size $30^{3}~{}\mathrm{fm}^{3}$, making Eq. (10A) unapplicable to other cases. For the other cases, Eq. (10C) describes the samples very well and can be considered as the uniform formulae to describe the energy shift with the the range of the potential as large as $10~{}\mathrm{fm}$. In comparison with the formula for the short-range potential, the power of $L$ in front of the exponential becomes larger. This enlighten us that the power of $L$ has a correlation with the range of the force. A theoretical formulation of the two-body quantization condition in the presence of a long-range force was proposed in  [6, 11, 7, 8, 9]. We expect that a comparable power-law dependence on $L$ can be obtained within the improved Lüscher formula.
As discussed above, the power of $L$ in front of the exponential is related to the range of the force. Accordingly, we study tendency of this power in the following. According to the results from machine learning approach, the formula can be written as a general form:

### v2

@@ Abstract
Finite-volume extrapolation is an important step for extracting physical observables from lattice calculations. However, it is a significant challenge for systems with long-range interactions. We employ symbolic regression to regress the finite-volume extrapolation formula for both short-range and long-range interactions. The regressed formula still holds the exponential form with a factor $L^{n}$ in front of it. The power decreases with the decreasing range of the force. When the range of the force becomes sufficiently small, the power converges to 1-1- 1, recovering the short-range formula as expected. Our work represents a significant advancement in leveraging machine learning to probe uncharted territories within particle physics.
@@
1. Introduction.   Lattice quantum chromodynamics, which discretizes space into a four-dimensional hypercubic lattice and performs the calculations in a given finite volume, is a foundational computational framework for studying strong interactions. It predicts not only the spectrum of conventional hadrons but also reveals the existence of exotic hadronic states. In particular, it has recently provided precise calculations for near-threshold states, taking the double charm tetraquark as an example. [1, 2, 3] To extract physical observables from lattice calculations, finite-volume extrapolation is essential. However, the presence of long-range interactions presents significant challenges, as they invalidate key assumptions underlying standard approaches, most notably the Lüscher formula. [4] When light particles mediate forces between scattering states, conventional methods may fail, leading to uncontrolled systematic uncertainties. Addressing these issues is crucial for reliable lattice analyses, particularly in systems where long-range effects play a dominant role, such as one-pion exchange in nucleon-nucleon scattering and the effects of light-meson exchange in the $D^{*}D$ system. [5] Recent theoretical studies [6, 7, 8, 9, 10] try to solve this problem by either modifying the Lüscher formula numerically, but without an explicit formula as simple as that for the short-range interaction, [11] or performing quantization conditions using a plane-wave basis. [12] Moreover, understanding finite-volume effects in two-body systems with long-range interactions is a necessary step towards addressing similar challenges in three-body systems.
@@
Fortunately, deriving fundamental mathematical expressions directly from empirical data [13] is an advantage of symbolic regression, [14] which has exhibited its capability in physics and scientific research. [15] Unlike traditional regression, it automatically discovers interpretable laws by exploring vast expression spaces. This has enabled the rediscovery of known physical laws and the formulation of novel equations for complex systems, [16] even when their underlying principles are elusive. Recent advances in symbolic regression enhance its effectiveness by incorporating physical constraints like dimensional consistency, thus improving search efficiency and model plausibility. For example, the AI Feynman algorithm has derived 100 equations from the Feynman Lectures on Physics. [17] Its ability to generate human-readable models accelerates discovery across various fields. [18, 19] For instance, in high-energy physics, it has revealed fundamental distributions, such as the Tsallis distribution in hadron transverse momentum analysis. [20] Moreover, it has provided analytical expressions for key low-energy observables, including the Higgs mass, muon anomalous magnetic moment, and dark matter relic density. [21] Reference [22] uses neural network to reproduce the numerical Lüscher’s formula to a high precision.
@@
As a result, we employ symbolic regression to probe the finite-volume extrapolation formula for long-range interactions. Before symbolic regression, we need to prepare samples for training symbolic regression. For these samples, we borrow the hadron lattice effective field theory (HLEFT), [23] which is analogous to the nuclear lattice effective field (NLEFT), a powerful computational framework for ab initio nuclear structure calculations. [24] The approach implements a systematic discretization of nucleon-nucleon interactions on a three-dimensional cubic lattice with periodic boundary conditions. These quantum many-body systems are subsequently solved using advanced numerical techniques including the Lanczos diagonalization method and auxiliary field Monte Carlo simulations. NLEFT has demonstrated remarkable success across diverse domains of nuclear physics, particularly in characterizing nuclear ground states [25, 26, 27, 28, 29, 30, 31] and excited states, [32, 33, 34, 35] elucidating nuclear intrinsic density distributions and clustering phenomena, [36, 37, 38, 39, 40] analyzing nucleus-nucleus scattering processes, [41, 42] and investigating both zero-temperature and finite-temperature nuclear matter properties. [43, 44, 45, 46] The framework has also been effectively extended to study hypernuclear systems. [47, 48, 49, 23] However, the presence of finite-volume effects remains a significant challenge in extending NLEFT applications to critical nuclear systems such as the neutron-rich ${}^{6}\mathrm{He}$ isotope, where the delicate balance of nuclear binding requires exceptional precision. This is another motivation of this work. According to the statement above, our workflow is presented in Fig. 1.
Figure 1: Workflow of using symbolic regression, LEFT, and formula.
@@
2. Samples Generation.   The HLEFT is used to generate the model-independent samples of two identical boson systems. These samples are fed to symbolic regression (SR), which is based on genetic algorithms (GA). [50, 51] GA allows the computer to require fewer prior assumptions and provides a more flexible search space to find final expressions. Similar to the evolutionary processes in nature, GA employs mechanisms such as mutations and crossovers to generate new solutions. In this work, we use the famous PySR, which is a multi-population evolutionary algorithm with multiple evolutionary processes performed asynchronously. It uses a classic evolutionary algorithm for each cycle and tournament selection for individual selection. The details of the setup in PySR can be found in the Supplementary Materials (SM).
@@
The challenge of finite volume extrapolation lies in long-range interactions, [6, 12, 7, 8, 9] especially when the range of force is comparable to the box size. Luckily, the extrapolation formula for the short-range potential can be obtained from Lüscher’s formula for $\kappa L\gg 1$[11] with $\kappa$ the binding momentum and $L$ the box size. Firstly, we regress the extrapolation formula for the short-range potential as a benchmark for the applicability of the PySR model. We work in a non-relativistic framework and consider two identical spinless particles with the same mass $m=1969~\mathrm{MeV}$.   [^1] The polarization of non-zero spin particles would modify the finite-volume extrapolation formula, which will not be discussed in this manuscript. The short-range potential in coordinate space reads as
$V(\bm{r})=-C_{0}\delta^{3}(\bm{r}),$ (1)
[…]
$H=\sum_{i=1}^{2}\frac{\bm{p_{i}}^{2}}{2m_{i}}+f(\bm{p}^{(\prime)}_{1},\bm{p}^{(\prime)}_{2})V(\bm{q}),$ (2)
@@
in which $m_{i}$ is the mass of the $i$th particle, $\bm{q}=\bm{p}-\bm{p}^{\prime}$ is the transferred momentum of the two particles and $\bm{p}$($\bm{p^{\prime}}$) represents relative incoming (outgoing) momentum between $\bm{p}_{1}$ ($\bm{p}^{\prime}_{1}$) and $\bm{p}_{2}$ ($\bm{p}^{\prime}_{2}$). The corresponding potential in momentum space $V(\bm{q})$ = $-C_{0}$ can be obtained by a Fourier transform from the expression in coordinate space.
@@
where $H_{L}$ and $E_{L}$ are the Hamiltonian and the binding energy in a cubic box $L^{3}$. The solution of Eq. (3) can be written explicitly in a single-particle basis $|\psi\rangle=|n_{1},\cdots,n_{N}\rangle$[44] where $n_{i}$ is an integer triplet specifying the lattice coordinates. The matrix exact diagonalization scheme uses the implicitly restarted Lanczos method to find the eigenvalues and eigenvectors [53] by SciPy [54] in Python. In order to study the finite volume extrapolation formula for short-range potential, we perform simulations on box sizes $L^{3}=10^{3},11^{3},\cdots,30^{3}~\mathrm{fm}^{3}$ cubic lattice with two identical particles. The potential strength $C_{0}$ is set to 1.5, 2.0, 2.5, and 3.0 $\mathrm{MeV}^{-2}$ to generate the samples. With this setup, we can extract the energy $E_{L}$ in finite volume by solving Eq. (3) for various cases. The details of the samples can be found in the SM.
The Yukawa potential, [6]
$V(\bm{r})=-C_{01}\delta^{3}(\bm{r})-C_{02}\frac{e^{-\mu r}}{r}$ (4)
[…]
@@
Here, both the short-range potential $V_{S}(\bm{q})$ = $-C_{01}$ and the long-range potential $V_{L}(\bm{q})\sim\frac{1}{\bm{q}^{2}+\mu^{2}}$ can be transformed into their coordinate forms by Fourier transformation. We perform simulations on box sizes $L^{3}=10^{3},11^{3},\cdots,30^{3}$ fm3 cubic lattices. We set the potential strength $C_{01}=C_{02}=0.03,~0.09,~0.12,\text{and}~0.21~\mathrm{MeV}^{-2}$ to obtain binding energies similar to those for only the short-range potential. The details can be found in the SM.
3. Symbolic Regression.  We perform symbolic regression on the generated samples that characterize finite-volume effects. The PySR model [55] samples the space of analytic expressions defined by the set of operators, input variables, and constant terms for minimization through genetic programming. The pool of operations includes addition, subtraction, multiplication, division, exponential, logarithm, square, etc. The input variable is box size $L$. Unlike the massive data used in deep learning, a few samples should be sufficient in the PySR model. The model evolution process employs several mechanisms, such as mutations and crossovers, to generate new expressions. The algorithm performs mutations over 50 iterations of 50 different population samples, with each population containing 35 individuals. There are two elements to measure the goodness of the output formula in the PySR model, i.e., loss and score. The value of loss is used to measure how well the output formula describes the samples, which is defined as mean square error,
@@
Here, $E_{\mathrm{PySR}}(L)$ is the output formula of PySR model, $E_{L}({L})$ is the energy at box size $L^{3}$ calculated from lattice effective field theory (LEFT). Score is used to estimate the form of the formula, which rewards minimal loss and penalizes the more complicated formula. It can be defined as
@@
Here, $C$ is the complexity  [^3] The upper limit of the $C$ value can be set larger to allow higher order contributions as that in analytic analysis  [11]. (its values can be found in the SM), which is defined as the total number of operations, variables, and constants used in a formula. The PySR model will consider both score, i.e. complexity and loss, to choose the best formula. The other parameters used in training process can be found in the SM.
@@
The training process of symbolic regression is very similar to biological population inheritance. In initialization, the PySR model randomly provides several kinds of evolutionary frameworks, which are different from each other and are used as a starting point for a given evolution. Mutations and crossovers occur throughout each iteration. During each iteration, formulae will be produced and the PySR model will choose the best formula for the next iteration according to the value of loss and score. The iterations will stop until the loss and complexity reach the specified values set at the beginning. During this process, the PySR model will choose the best formula across all iterations.
@@
Figure 2: The results of fitting the formula to samples. There are four cases of (a) $C_{0}=1.5$ MeV-2, (b) $C_{0}=2.0$ MeV-2, (c) $C_{0}=2.5$ MeV-2 and (d) $C_{0}$ = 3.0 MeV-2. Equations (a), (b) and (c) correspond to Eqs. (8a), (8b) and (8c), respectively. The box size $L$ ranges from 10 to 30 fm. The purple region and the yellow region indicate the fitted and predicted areas by Eqs. (a), (b) and (c), respectively. Each graph has its own sub-graph to show greater detail of the lines between the fitted and predicted areas.
@@
4. Results.  Based on the PySR model, we can perform a symbolic regression for the samples of the short-range attractive potential to find an expression that uniformly describes the finite-volume extrapolation. In the four cases of $C_{0}$ values, three formulae are accepted by the PySR model,  [^4] For the case of $C_{0}=1.5~\mathrm{MeV}^{-2}$, PySR model gives the formula $E_{L}=C^{\prime}_{1}+C^{\prime}_{2}/L+C^{\prime}_{3}/L^{2}$. For the purpose of uniformity of the formula, when the ratio $C^{\prime}_{2}/C_{3}^{\prime}$ is small enough, the formula is approximated as $E_{L}=C_{1}+C_{2}e^{-C_{3}L}/L^{2}$. as follows:
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L^{2},$ (8a)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L,$ (8b)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}.$ (8c)
@@
In Fig. 2, we show the results of the three formulas fitting to the samples with the short-range potential, generated from LEFT. The purple area is the fitting region for the formulae derived from the PySR model. The yellow area is the predicted one. For the shallow bound state, e.g., Fig. 2(a), equation (8a) is the best choice. However, the volume size is not big enough to see the convergent behavior, making the regressed formula not universal. When the binding energy increases to a few MeV, i.e., Figs. 2(b) and  2(c), equation (8b) is better than Eq. (8a). When the binding energy further increases to more than $10~\mathrm{MeV}$, e.g., Fig. 2(d), all three formulae are good choices because their loss values are much smaller than those of the other sub-figures of Fig. 2. As a result, we conclude that Eq. (8b) is the best regressed uniform formula to describe the samples with a few MeV binding energy.
@@
The current well-established theoretical works [4, 56, 57, 58] are based on the Lüscher formula [11] for the short-range potential with $\kappa L\gg 1$, where $\kappa$ is the binding momentum of the two-body system. The energy shift of a shallow two-body bound state of identical bosons in finite volume is
@@
where $C^{\prime}$ is the free parameter. This formula is exactly the form of Eq. (8b) regressed by PySR. In addition, the fitted values of the parameter $C_{3}$ in Eq. (8b) for the last three cases of Fig. 2 are equal to the binding momentum $\kappa$ numerically. This also demonstrates that we have successfully regressed the finite-volume extrapolation formula for the short-range potential.
[…]
For the long-range potential, the strengths of the attractive potentials are set to obtain similar binding energies for the short-range potential. The range of the force parameter $\mu$ is set as $\mu=20~\mathrm{MeV}$ for these four $C_{0}$ cases. The other values of the force parameter and their effect on the finite-volume extrapolation will be discussed afterwards. The following three formulae  [^5] For the case of $C_{01}=C_{02}=$ 0.03 MeV-2, PySR model gets the formula $E_{L}=C^{\prime}_{1}+C^{\prime}_{2}/L+C^{\prime}_{3}L$. For the purpose of a uniform form of the formula, when the ratio $C^{\prime}_{3}/C^{\prime}_{2}$ is small enough, the formula is approximated as $E_{L}=C_{1}+C_{2}e^{-C_{3}L}/L$.
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}/L,$ (10a)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L},$ (10b)
$E_{L}$ $=C_{1}+C_{2}e^{-C_{3}L}L.$ (10c)
are accepted by PySR model. Figure 3 shows the fitting results of the above three formulae to the energy with the long-range interaction. For the case with binding energy smaller than $1~\mathrm{MeV}$, i.e., figure 3(a), equation (10a) is preferred. However, the binding energy is too small to make the energy convergent in box size $30^{3}~\mathrm{fm}^{3}$, making Eq. (10a) inapplicable to other cases. For the other cases, equation (10c) describes the samples very well and can be considered as a uniform formula to describe the energy shift with the range of the potential as large as $10~\mathrm{fm}$. In comparison with the formula for the short-range potential, the power of $L$ in front of the exponential becomes larger. This enlightens us that the power of $L$ has a correlation with the range of the force. A theoretical formulation of the two-body quantization condition in the presence of a long-range force was proposed in Refs. [6, 12, 7, 8, 9, 10]. We expect that a comparable power-law dependence on $L$ can be obtained within the improved Lüscher formula.
As discussed above, the power of $L$ in front of the exponential is related to the range of the force. Accordingly, we study the tendency of this power in the following. According to the results from the machine learning approach, the formula can be written as a general form,